OpenStudy (anonymous):

if y^8=4 and y^7=3/x, what is the value of y in terms of x?

OpenStudy (anonymous):

not sure what " what is the value of y in terms of x?" means

OpenStudy (anonymous):

if \(y^8=4\) then \(y=\sqrt[8]{4}\) or \(y=-\sqrt[8]{4}\)

OpenStudy (anonymous):

only a choice of two number

OpenStudy (anonymous):

we have to find y like these are the answer choices: a. 4x/3 b. 3x/4 c. 4/x d. x/4 e. 12/x

OpenStudy (anonymous):

there must be a typo somewhere in the problem

OpenStudy (anonymous):

why? what do u mean?

OpenStudy (anonymous):

if \(y^7=\frac{3}{x}\) then \[y=\sqrt[7]{\frac{3}{x}}\]

OpenStudy (anonymous):

im not really sure how to do this but i checked the answer and it says that the answer is 4x/3

OpenStudy (anonymous):

maybe we can try this \[y^8=4, y^7=\frac{3}{x}\] and \[y^8=y^7\times y\] so \[y^8=y^7\times y\implies 4=y^7\times y\implies 4=\frac{x}{3}y\]

OpenStudy (anonymous):

typo there sorry

OpenStudy (anonymous):

\[y^8=y^7\times y\implies 4=y^7\times y\implies 4=\frac{3}{x}y\]

OpenStudy (anonymous):

its ok :)

OpenStudy (anonymous):

then solve for \(y\)

OpenStudy (anonymous):

ok :) i got it thanks a lot :)

OpenStudy (anonymous):

yw